OpenStudy (anonymous):

x(x+4)-5(x+4) can someone factor

OpenStudy (anonymous):

is this factorable

OpenStudy (radar):

(x=4)(x-5)

OpenStudy (radar):

(x+4)(x-5)

OpenStudy (anonymous):

(X-5)(X+4) this is how you do it... this is how i got this.. when you simplify you get x^2-x-20. so... -4*5=-20 and -5+4=-1. so you have to refer back to your simplified equation. :) hope i make sense

OpenStudy (anonymous):

You don't really need to do all that math... The equation is just distributed out a little bit. It started with (x-5)(x+4) and they just distributed the first one. They could have done the same the other way and the problem would have been x(x-5)+4(x-5)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

the reason that i showed you that way is when you get to more complex and 3 variables to work with you are giong to have to know what you need too. it makes it less complicated. i hope this is enough advise:):)